9. The following table shows the peak wind gust (X) and the number of damaged homes (Y) (in thousands) in seven cities in a tPlease No graphs. I need help solving the problem with the computations
Answer #1

I am going to provide you a method to estimate a and b of the given model.

Given model is y = ax​​​​​b​ .

Taking logarithm base e on both sides we get

log_e y = log_e(ax^b)\\log_ey =log_ea + log_ex^b\\ log_ey =log_ea + b\cdot log_ex \\ z= c + b\cdot u

Where z=logey, c = log​​​​e​​​a, u = log​​​​e x

Calculate all the values of corresponding z and u.

Now this gives an equation with two unknowns and 7 data points (z,u).

We can solve this equation with the given data, to find the value of c and b.

estimate of c is given by mean of the z values. Then by taking antilog of c we get a, i.e. a = e​​​​​​c

The estimate of b is given by \frac{\sum_{i=1}^{7}z_i \cdot u_i}{\sum_{i=1}^{7} (u_i)^2} .

To show that it is a good fit, compute the MSE, i.e. \frac{\sum_{i=1}^{7} (z_i - \hat{z_i})^2}{7}

Where the \hat{z_i} are obtained from the estimated equation putting in the values of known u values.

Thank you. Hope this helps you with the computation.